# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v2

@@ 2.4.2 Performance Metrics
(a) Segmentation by 1D-Justo-LiuNet.
(b) Segmentation by 2D-CUNet++ Reduced.
Figure 6: Segmentation results for a HYPSO-1 capture over the Namib desert in Namibia, Africa, close to the Gobabeb Namib Research Institute for dry land research. The approximate coordinates for this location are 23°19’06.0”S latitude and 14°44’21.8”E longitude. The grayscale image in (a) and (b) corresponds to the same L1b channel with a wavelength $\lambda$ of 699.61 nm, close to NIR spectrum. The term SAMPLES denotes the swath width captured across the direction of satellite travel, while LINES indicates the number of scan frames collected along the satellite’s path. As stated in [5, 29], HYPSO-1’s spatial resolution in the across- and along-track directions is approximately 100 m x 600 m.
3 Results
We conduct model training using 112 channels of both raw sensor data and L1b calibrated radiance, and we discuss the segmentation of raw and L1b radiance in Section 4. For L1b radiance, we present the outcomes of the segmentation models in Table II, where we utilize bold font to accentuate our models. The table displays the models’ performance for segmentation, with the primary metric being the average accuracy across the sea, land, and cloud classes. Additionally, to assess the sea-land-cloud segmentation performance for the context of the proposed on-board ranking system, we also include the average Spearman’s rank coefficient. In particular, the average Spearman’s coefficient across the respective rankings based on sea, land, and cloud coverage levels is presented in Table II. Moreover, we also include information about the network’s parameter count to offer insights of the potential memory demands for in-orbit deployment. In addition, we provide the average inference time per HS image measured in our testing setup on ground, on the NVIDIA RTX A4000 GPU (for DL), and on the 13th Gen Intel(R) Core (TM) i9-13900K CPU (for ML). From the numbers presented in Table II, we sort the models in Table III based on their respective metrics. In Table III, we first sort the models solely based on their average accuracy, from highest to lowest, and subsequently, we sort them in the following columns, first by the Spearman’s coefficient (highest to lowest), then by the number of parameters (lowest to highest), and finally by inference time (shorter to longer). Therefore, the first rows in the table represent the top models in the test set for each respective metric. However, while average accuracy offers a general overview on segmentation performance across the sea, land, and cloud classes, it may obscure poor model performance in specific classes. To address this, we offer additional insights into individual class performance by presenting their respective binary false detection ratios. Namely, in Figs. 3, 4, and 5, we provide a comparative visual representation of false alarm ratios for the binary segmentation of each distinct class. We note that the top-performing models get positioned in the lower-left corners, indicating the optimal trade-off balance between low false positives and low false negatives. Within the context of the example case presented in Fig. 2, the graphs illustrate the false alarm ratio on the Y-axis corresponding to the loss (or degradation) of valuable data, while the X-axis pertains to a false detection ratio that reflects the inefficient utilization of the downlink channel. In relation to this trade-off, we stress that while minimizing the loss of valuable data remains crucial, we also highlight the importance of maintaining an acceptable downlink efficiency.
Finally, our experiments conclude by training again the models using the 3-channel dataset instead. We employ the dataset to assess the impact of channel reduction on the segmentation models’ performance, parameter count, and inference times, as shown in Table IV. We note that the table omits the 1D models due to the 3-channel dataset’s restricted spectral information, as mentioned earlier in Section 2.3.4. Moreover, the empirical results are further confirmed by training once again the models with a set of three channels different from the ones in Section 2.2/A. The new results align with the previous findings and for simplicity, we include the new figures in the supplementary material.
Finally, as previously noted, the dataset predominantly contains instances of sea and land, with clouds representing the minority class. Consequently, the test set also exhibits a lower proportion of cloud data (14.13%), in contrast to sea (39.01%) and land (46.85%). We recognize this as a limitation that demands future refinement. Therefore, to mitigate the impact of the class imbalance during the testing phase, we conduct additional experiments. Namely, we perform inference on an additional on-ground deployment set, with 30 HS images, specifically designed to test the model’s generalization on e.g. cloudy and overexposed images to further support the findings of our work, and we provide the complete results in the supplementary material. As an example, Fig. 6 provides a comparison of the segmentation outcomes of the 1D-Justo-LiuNet and the 2D-CUNet++ Reduced for a sample image from the deployment set.
4 Discussion
4.1 Results Analysis for Models Trained on 112 Channels
We first analyze the results for the models trained on 112 channels (Tables II and LABEL:Table:models_sorted_according_to_performance). Our objective is to investigate how the utilization of the imaging spatial context by 2D models affects the sea-land-cloud segmentation across the test images. We assess the models’ strengths and limitations in terms of performance, spanning from the average accuracy and the binary false alarm ratios to the Spearman’s correlation coefficient for the ranking system. We finalise the analysis by discussing the model’s parameter count and inference time, which are relevant metrics for on-board deployment. Prior to the analysis of the results, we intuitively formulate a hypothesis next, independent of the empirical findings. The hypothesis asserts that it is reasonable to anticipate better performance for 2D models compared to 1D models. This expectation arises because 2D models, unlike 1D models relying solely on spectral context, analyse a richer context comprising both spectral and spatial information. We next proceed to elaborate whether the empirical results substantiate such hypothesis.
4.1.1 Segmentation Performance (I): Average Accuracy
When examining the segmentation performance solely based on the average accuracy across sea, land, and cloud classes, we observe that our proposed model, 1D-Justo-LiuNet, achieves the highest accuracy of 0.93 in Table II. The result of a 1D model outperforming 2D models is not an exception, as four out of five models ranking among the top performers in accuracy in Table LABEL:Table:models_sorted_according_to_performance are consistently 1D models. The only exception in this top five ranking is our 2D model, 2D-Justo-UNet-Simple, which achieves an accuracy of 0.92, closely approaching the performance of the 1D-Justo-LiuNet. Therefore, the findings lead to the conclusion that the initially expected hypothesis suggesting that 2D models would outperform 1D models does not hold valid in our experiments, as the 2D models exhibit worse performance. This phenomenon can be attributed to the fact that 1D models, unlike 2D using the spatial context, are not affected by the modest spatial resolution of HYPSO-1’s images while not suffering from the spatial variations that may be practically irrelevant and misleading for pixel-level segmentation. Instead, 1D models dedicate their efforts to analyzing exclusively the spectral context, which is, certainly, a powerful domain in HS imagery. Additionally, Table LABEL:Table:models_sorted_according_to_performance demonstrates that our proposed models (1D-Justo-LiuNet, 2D-Justo-UNet-Simple, 1D-Justo-HuNet, and 1D-Justo-LucasCNN) outperform, in average accuracy, the state-of-the-art models under analysis, when trained on the HYPSO-1 dataset for sea-land-cloud segmentation. Indeed, we surpass models such as the 2D-NU-Net-mod, 2D-NU-Net, 2D-UNet FAUBAI, and 2D-UNet FAUBAI Reduced previously employed in the field of cloud segmentation for satellite remote sensing imagery. Finally, we highlight that the average accuracy of 1D-Justo-LiuNet (0.93), 1D-Justo-HuNet (0.91), and 1D-Justo-LucasCNN (0.91) also surpass the literature’s average accuracy for soil texture [45]. As a matter of fact, the state of the art achieves significantly lower average accuracy of 0.59, 0.61, and 0.56 for Liuetal, Huetal, and LucasCNN, respectively.
4.1.2 Segmentation Performance (II): The Effect of L1b Radiance Calibration
The models in Table II are trained on L1b calibrated radiance data. L1b radiance is one of the standard data formats in HS imagery, which not only allows satellite comparisons but also removes sensor-specific characteristics and biases. Next, we empirically examine how L1b radiance calibration impacts segmentation performance. In the supplementary materials of this article, we provide the average accuracy, along with other performance metrics, for the models trained on raw sensor data. This supplementary information is provided to facilitate comparisons with the models trained on L1b radiance, as presented in Table II, in support of the upcoming analysis. To begin, for the highest-performing DL models in terms of average accuracy shown in Table III (1D-CNNs), the models exhibit slight superior performance with L1b radiance compared to training them with raw data. A similar trend occurs for the 2D models ranked just below the 1D models, which also demonstrate enhanced performance with radiance data versus raw data. Conversely, the models at the lower end of the accuracy ranking tend to perform better with raw data than radiance calibration. Concentrating only on the superior models in Table III, which hold more significance as they are the top performers, we conclude that L1b radiance calibration tends to positively impact and improve segmentation performance. We attribute the improvement to the correction of errors and biases during L1b radiance calibration, allowing top-performing models to focus on inherent data patterns rather than overfitting on sensor-induced characteristics, leading to improved generalization performance. Nonetheless, we strongly recommend conducting comparative testing on both calibrated and raw sensor data when training the models on datasets distinct from this paper to verify whether calibrated data, such as radiance or reflectance, continues to enhance model performance.
@@ 4.1.2 Segmentation Performance (II): The Effect of L1b Radiance Calibration
4.1.3 Segmentation Performance (III): False Detection Ratios, Data Loss, and Downlink Inefficiency
@@
For a more in-depth assessment of the segmentation performance for each individual class, we scrutinize which models are positioned in the lower-left corner in the scatter plots in Figs. 3, 4, and 5. The lower-left region ensures the most optimal false detection ratios, and for the example ranking system in Fig. 2, the region achieves minimal loss (or degradation) of valuable data, while also minimizing the inefficient utilization of the satellite downlink. As previously explained, lower data loss (in the Y-axis) takes precedence, while the downlink inefficiency (in the X-axis) is of lower priority since it does not result in any loss of valuable data. However, as discussed in Section 2.4, we also aim for a reasonably balanced relationship between false positives and false negatives to prevent scenarios where a model achieves a convenient minimal data loss, at the expense of extremely deteriorating the downlink efficiency. Next, we comparatively analyze the top-performing models located in the lower-left corners of Figs. 3-5.
@@
First, we assess the models only in terms of data loss. In Fig. 3 for cloud segmentation, models such as the 2D-Justo-UNet-Simple excels in minimizing the loss of data, approaching nearly no data loss (ideal case). This signifies that the 2D-Justo-UNet-Simple outperforms by not misclassifying valuable non-cloud data (sea or land) as clouds, even if it means some cloudy data remains undetected, thus reducing the downlink efficiency. With respect to sea segmentation (Fig. 4), the 2D-UNet-FAUBAI-reduced stands out for its minimal loss of sea pixels with FNR=0.01, i.e., there is only a 1% proportion of sea pixels that remain undetected, resulting in their loss. This outperforms other models such as the 1D-Justo-LiuNet with 5% loss. Finally, in Fig. 5 for land segmentation, we observe a remarkable similarity in false alarm ratios in the lower-left optimal region, resulting in a slight overlap. Therefore, we next clarify that the optimal models, when sorted in ascending order of FNR (data loss), are as follows: 0.02 (2D-Justo-UNet-Simple), 0.03 (1D-Justo-LiuNet), 0.04 (1D-Justo-HuNet), and 0.06 (2D-CUNet++ Reduced). Consequently, the 2D-Justo-UNet-Simple is the most optimal at minimizing the loss of land data, achieving only a 2% ratio of undetected land pixels.
@@
Second, to jointly maintain a reasonable downlink efficiency while achieving low data loss, we study the models taking into account not only the data loss as in the previous analysis, but also the trade-off between both loss of data and downlink efficiency objectives. We could consider the use of Pareto fronts, which are typically applied in multi-objective optimization to address trade-offs between inversely proportional variables. Nevertheless, for simplicity, we analyze the trade-off using the Euclidean distance metric that gives equal weight to the two objectives. Specifically, we calculate the distance of each model from the origin point, representing the ideal scenario with zero false alarms. First, for cloud segmentation (Fig. 3), the models closest to the ideal performance have (FPR, FNR) values of (0.02, 0.15) for 1D-Justo-LiuNet and (0.00, 0.17) for 2D-Justo-UNet-Simple, resulting in approximate distances of 0.151 and 0.170, respectively. In brief, 1D-Justo-LiuNet exhibits the shortest distance to the ideal model. Second, the top models for sea segmentation (Fig. 4), with (FNR, FPR) of (0.08, 0.01) for 1D-Justo-LucasCNN, (0.05, 0.02) for 1D-Justo-LiuNet, and (0.04, 0.03) for 2D-UNet-FAUBAI, have approximate distances of 0.081, 0.054, and 0.050, respectively. This suggests a marginally better performance of 2D-UNet-FAUBAI, with a difference of 0.004 in distance compared to the 1D-Justo-LiuNet. However, Table II notes that the 2D-UNet-FAUBAI has over 26 million parameters, whereas our network, 1D-Justo-LiuNet, has only 4,563 parameters, which is more suitable for on-board deployment. Considering the nearly identical performance and the large parameter difference against FAUBAI’s network, 1D-Justo-LiuNet remains the top model also for sea segmentation. Finally, for land segmentation (Fig. 5), the optimal models are of (FNR, FPR) = (0.02, 0.06) for 2D-Justo-UNet-Simple, (0.03, 0.04) for 1D-Justo-LiuNet, (0.04, 0.04) for 1D-Justo-HuNet, and (0.06, 0.04) for 2D-CUNet++ Reduced, with approximate distances of 0.063, 0.050, 0.057, and 0.072, respectively. In summary, 1D-Justo-LiuNet outperforms the other models for land segmentation.

### v3

@@ 7 FastViT-S12 2D-CUNet Reduced 1D-Justo-LucasCNN 1D-ML/LDA
As previously noted, the HYPSO-1 dataset from [34] predominantly contains instances of sea and land, with clouds representing the minority class. Consequently, the test set also exhibits a lower proportion of cloud data (14.13%), in contrast to sea (39.01%) and land (46.85%). We recognize this as a limitation that may demand future refinement. To mitigate the impact of class imbalance during the testing phase, we conduct additional experiments. Namely, aligning with the segmentation literature for satellite imagery [13], we perform inference on an additional on-ground deployment set, consisting of 30 HS images that are unseen by the model, to further test the model’s generalization ability for HYPSO-1. We compare the performance of 1D models with 2D models on this deployment set, including cloudy and overexposed images. As an example, Figs. 6-17 depict how 1D models generalize compared with 2D models on a sample image from the deployment set. This image was captured by HYPSO-1 on 20 October 2022, over the Namib desert in Namibia, Africa, near the Gobabeb Namib Research Institute for dry land research [34]. The approximate coordinates of this location are 23°19’06.0”S latitude and 14°44’21.8”E longitude. For each figure, we present, from left to right: the RGB composite of the image captured by the satellite; the segmentation results; the model’s confidence for each pixel-level prediction measured in terms of probability, with the highest confidence indicated by probabilities near the ideal value of 1.0 (100%), representing no uncertainty; and finally, a pie chart showing coverage levels for the sea, land, and cloud categories to sort the images in the ranking system. In the figures, the term SAMPLES denotes the swath width captured across the direction of satellite travel, while LINES indicates the number of scan frames collected along the satellite’s path. Finally, we note that the supplementary material includes further results for the remaining images in the deployment set, with inference run by 1D-Justo-LiuNet and 2D-CUNet++ Reduced.
Figure 6: Segmentation by 1D-Justo-LiuNet.
Figure 7: Segmentation by 1D-Justo-HuNet.
Figure 8: Segmentation by 1D-Justo-LucasCNN.
Figure 9: Segmentation by 2D-NU-Net-mod.
Figure 10: Segmentation by 2D-NU-Net.
Figure 11: Segmentation by 2D-CUNet.
Figure 12: Segmentation by 2D-CUNet++.
Figure 13: Segmentation by 2D-CUNet Reduced.
Figure 14: Segmentation by 2D-CUNet++ Reduced.
Figure 15: Segmentation by 2D-UNet FAUBAI.
Figure 16: Segmentation by 2D-UNet FAUBAI Reduced.
Figure 17: Segmentation by 2D-Justo-UNet-Simple.
III-B Generalization Results
Despite our previous experiments on HYPSO-1 data used L1b calibration, a universal satellite data format, we next conduct additional tests on other datasets to evaluate generalization and validate our upcoming conclusions. Table V presents our results on NASA’s EO-1 HS dataset for segmenting image pixels under three classes: no cloud, thin cloud, and thick cloud. As we will discuss, the 1D-CNN 1D-Justo-LiuNet outperforms the state of the art on HS data. To verify this, we compare the top-performing 1D-Justo-LiuNet with a high-performing 2D-CNN in terms of accuracy and Dice coefficient. Finally, we also conduct experiments on the RGB Satellite Imagery of Dubai (SID) dataset. Table VI presents the results for pixel-level segmentation across six classes: buildings, roads, land, vegetation, water, and other. The results include all 19 models previously tested for HYPSO-1.
TABLE V: Metrics comparing top-performing 1D-CNN, 1D-Justo-LiuNet, with a high-performing 2D-CNN, tested on NASA’s EO-1 hyperspectral dataset for generalization.
Channels Model Accuracy F1 Score
98 1D-Justo-LiuNet 0.82 0.97
nnU-net 2D 0.81 0.87
2D-Justo-UNet-Simple 0.70 0.81
6 1D-Justo-LiuNet 0.74 0.85
nnU-net 2D 0.80 0.85
2D-Justo-UNet-Simple 0.72 0.86
1 1D-Justo-LiuNet 0.64 0.76
nnU-net 2D 0.77 0.86
2D-Justo-UNet-Simple 0.57 0.88
TABLE VI: Metrics comparing 1D ML, 1D-CNNs, 2D-CNNs and ViTs models when tested on RGB Satellite Imagery of Dubai (SID) dataset for generalization.
Type Model Accuracy Dice Coefficient
1D 1D-ML/SGD 0.54 0.47
1D-ML/NB 0.59 0.50
1D-ML/LDA 0.64 0.53
1D-ML/QDA 0.65 0.54
1D-Justo-LiuNet 0.63 0.51
1D-Justo-HuNet 0.67 0.58
1D-Justo-LucasCNN 0.68 0.58
2D 2D-NU-Net-mod 0.75 0.65
2D-NU-Net 0.78 0.69
2D-CUNet 0.77 0.68
2D-CUNet++ 0.77 0.67
2D-CUNet Reduced 0.75 0.64
2D-CUNet++ Reduced 0.75 0.65
2D-UNet FAUBAI 0.82 0.70
2D-UNet FAUBAI Reduced 0.83 0.72
2D-Justo-UNet-Simple 0.76 0.66
FastViT-T8 0.82 0.71
FastViT-S12 0.82 0.70
FastViT-MA36 0.83 0.72
IV Discussion
IV-A Discussion of Main Results
IV-A1 Segmentation Performance: Average Accuracy
When evaluating the segmentation performance based on the average accuracy across sea, land, and cloud classes, we observe that the proposed model, 1D-Justo-LiuNet, achieves the highest accuracy at 0.93, as shown in Table II. Our model surpasses the remaining 18 state-of-the-art models, including 2D-CNNs, and outperforms all FastViT configurations, including its lightest configuration, FastViT-T8, and its heaviest version, FastViT-MA36. The result of a 1D model outperforming 2D models is not an exception, as Table III confirms that four out of the top five models in terms of accuracy are 1D models. The only exception is our 2D model, 2D-Justo-UNet-Simple, which achieves an accuracy of 0.92, closely approaching the performance of the 1D-Justo-LiuNet. These results can be attributed to the simplicity of 1D-CNNs, which extract features solely from the spectral domain. With fewer parameters and lower complexity, the risk of overfitting is reduced. In contrast, 2D models incorporate spatial context, significantly increasing the number of parameters and complexity, which raises the likelihood of overfitting, especially when training data is limited, making them more prone to learning spurious features. In short, 1D-CNNs achieve superior results in HS segmentation due to their simplicity, aligning with the principle of Occam’s razor, which favors the simplest model as the best solution. We will demonstrate later that this conclusion also generalizes to other hyperspectral datasets.
IV-A2 Segmentation Performance: Priority Rankings for Downlink (Case Example)
In this work, we simulate on ground a ranking system for future on-board deployment. From the labels in the 5 test images, we calculate the proportions of sea, land, and cloud classes. We emulate a downlink prioritization system based on cloud coverage, sorting from lowest cloud coverage (higher priority) to highest (lower priority). This serves as the ground-truth ranking based on cloud coverage. Next, a given model segments the images, and from these segments, we calculate the predicted class proportions. Using these proportions, we generate a predicted ranking based on cloud coverage, starting from lower to higher. Ideally, the predicted ranking would match the ground-truth ranking. To assess this, we compare the two rankings using Spearman’s correlation coefficient. This allows us to see how the given model performs for a case application where the segments serve to guide data downlink to Earth, optimizing satellite operations and automation.
Table II shows that 1D-Justo-LiuNet, along with 1D-Justo-HuNet, outperforms the state of the art by achieving the highest possible performance, with a Spearman’s coefficient of 1.00, resulting in an error-free ranking. Consistent with the accuracy results, Table III further confirms that four out of the top five models, now in terms of Spearman’s coefficient, are 1D models. As we will discuss in the next section, the table also highlights that the parameter count for the lightest FastViT-T8 is in the order of millions, which is significantly less suitable for in-orbit deployment compared with the few thousand parameters required by 1D-Justo-LiuNet and 1D-Justo-HuNet. Due to the high memory usage for storing parameters, FastViTs are not a suitable option for in-orbit deployment. Therefore, we exclude FastViTs from the ranking experiment. Moreover, we emphasize that FastViTs cannot surpass the ideal Spearman’s coefficient of 1.00, already achieved by 1D-Justo-LiuNet and 1D-Justo-HuNet.
IV-A3 Parameter Count
Table II focuses on the parameter count for deep models, as these typically have larger number of parameters compared with conventional ML algorithms. Parameter count is crucial for inference at edge due to the limited memory resources available on board. Tables II and III show that 1D-Justo-LiuNet, which achieves top segmentation results, also outperforms all other state-of-the-art deep models in terms of parameter count, making it the lightest deep model. Thus, 1D-Justo-LiuNet is a highly compressed deep model that not only maintains but exceeds the segmentation performance of both conventional and the latest state-of-the-art models. We conclude that our network is the most suitable option for future deployment aboard HYPSO-1. To this extent, as an illustrative case, HYPSO-1 uses a Zynq-7030 System-on-Chip (SoC) with a dual-core ARM Cortex-A9 CPU and a Kintex-7 FPGA for on-board data processing [10]. The dual-core CPU, limited by power constraints, has only 32 KB of L1 data cache for fast memory access [50], theoretically allowing up to 8,000 model parameters to be stored (assuming 4-byte encoding). This illustrates the limited memory resources available on board. In our case, 1D-Justo-LiuNet requires only a few kilobytes for parameter storage, whereas even the lightest FastViT-T8, with millions of parameters and a higher risk of overfitting, demands considerably larger memory - in the order of megabytes. This makes FastViTs more challenging for in-orbit deployment.
IV-A4 Inference Time
Table II shows that 2D-CNNs take longer inference times compared with 1D-CNNs and ViTs. For this analysis, we focus on the lightest models: 1D-Justo-LiuNet, 1D-Justo-HuNet, and ViTs such as FastViT-T8 and FastViT-S12. Although FastViT-T8 has a faster inference time of 132 ms, 1D-Justo-LiuNet is only 34 ms slower at 166 ms. However, 1D-Justo-LiuNet achieves significantly higher accuracy at 0.93, compared with FastViT-T8’s 0.82. Considering this accuracy boost and the fact that 1D-Justo-LiuNet has over 1,000 times fewer parameters, we conclude that our model turns out as the best-performing model for in-orbit deployment. Furthermore, FastViT-T8, despite its speed, is not the fastest model. Our model 1D-Justo-HuNet achieves the most optimal inference time at just 69 ms, which is 63 ms faster than FastViT-T8. With an accuracy of 0.91 and a slightly higher parameter count than 1D-Justo-LiuNet, 1D-Justo-HuNet still outperforms all FastViTs in terms of accuracy and parameter counts, while also surpassing their inference time. Indeed, 1D-Justo-HuNet surpasses all the remaining 18 state-of-the-art models. Therefore, it is also a very feasible alternative for in-orbit deployment.
@@ IV-A4 Inference Time
IV-A5 Class Segmentation Performance: False Detection Ratios, Data Loss, and Downlink Inefficiency
@@
For a more in-depth assessment of the segmentation performance at the class level, we scrutinize which models are positioned in the lower-left corner in the scatter plots in Figs. 3-5. Given our earlier discussion on FastViTs and their large number of parameters for in-orbit deployment, the figures instead focus on lighter ML and DL models. The lower-left region ensures the most optimal false detection ratios, and within the example ranking system, the region achieves minimal loss (or degradation) of valuable data, while also minimizing the inefficient utilization of the satellite downlink. As mentioned, lower data loss (in the Y-axis) takes precedence, while the downlink inefficiency (in the X-axis) is of lower priority since it does not result in any loss of valuable data. However, we still balance a reasonable relation between false positives and false negatives to prevent scenarios where a model achieves minimal data loss at the cost of a highly inefficient use of downlink resources.
@@
Next, we comparatively analyze the top-performing models located in the lower-left corners of Figs. 3-5. To begin, we assess the models only in terms of data loss before we include downlink efficiency. In Fig. 3 for cloud segmentation, models such as 2D-Justo-UNet-Simple excels in minimizing the loss of data, with nearly no data loss (ideal case). It means that the 2D-Justo-UNet-Simple outperforms by not misclassifying valuable non-cloud data (sea or land) as clouds, even if it means some cloudy data remains undetected, thus reducing the downlink efficiency. With respect to sea segmentation (Fig. 4), the reduced 2D-UNet FAUBAI stands out for its minimal loss of sea pixels with FNR=0.01, i.e., there is only a 1% proportion of sea pixels that remain undetected, resulting in their loss. This outperforms other models such as the 1D-Justo-LiuNet with 5% loss. Finally, in Fig. 5 for land segmentation, we observe a remarkable similarity in false alarm ratios in the lower-left optimal region, resulting in a slight overlap. Therefore, we next clarify that the optimal models, when sorted in ascending order of FNR (data loss), are as follows: 0.02 (2D-Justo-UNet-Simple), 0.03 (1D-Justo-LiuNet), 0.04 (1D-Justo-HuNet), and 0.06 (2D-CUNet++ Reduced). Consequently, the 2D-Justo-UNet-Simple is the most optimal at minimizing the loss of land data, achieving only a 2% ratio of undetected land pixels. However, although data loss is certainly critical, this analysis does not consider the balance between data loss and downlink efficiency. Therefore, to jointly maintain a reasonable efficiency while achieving low data loss, we next study the models taking into account this trade-off.
@@
We could consider the use of Pareto fronts, which are typically applied in multi-objective optimization to address trade-offs between inversely proportional variables, as is the case here. Nevertheless, for simplicity, we analyze the trade-off using the Euclidean distance, which gives equal weight to the two objectives. Namely, we calculate the distance of each model from the origin point, representing the ideal scenario with zero false alarms. First, for cloud segmentation (Fig. 3), the models closest to the ideal false-free scenario have (FPR, FNR) values of (0.02, 0.15) for 1D-Justo-LiuNet and (0.00, 0.17) for 2D-Justo-UNet-Simple, resulting in approximate Euclidean distances of 0.151 and 0.170, respectively. Therefore, 1D-Justo-LiuNet exhibits the shortest distance to the ideal scenario. Second, the top models for sea segmentation (Fig. 4) - with (FNR, FPR) of (0.08, 0.01) for 1D-Justo-LucasCNN, (0.05, 0.02) for 1D-Justo-LiuNet, and (0.04, 0.03) for 2D-UNet-FAUBAI - have approximate distances of 0.081, 0.054, and 0.050, respectively. This suggests a better performance for 2D-UNet-FAUBAI, with a marginal difference of 0.004 in distance compared with the 1D-Justo-LiuNet. However, Table II notes that the 2D-UNet-FAUBAI has over 26 million parameters, whereas our network, 1D-Justo-LiuNet, has only 4,563 parameters, which is clearly more suitable for on-board deployment. Balancing the nearly identical performance and the large parameter difference against FAUBAI’s network, 1D-Justo-LiuNet remains the top model for sea segmentation as well. Finally, for land segmentation (Fig. 5), the optimal models are of (FNR, FPR) = (0.02, 0.06) for 2D-Justo-UNet-Simple, (0.03, 0.04) for 1D-Justo-LiuNet, (0.04, 0.04) for 1D-Justo-HuNet, and (0.06, 0.04) for 2D-CUNet++ Reduced, with resulting distances of 0.063, 0.050, 0.057, and 0.072, respectively. Therefore, 1D-Justo-LiuNet outperforms the other models for land segmentation.